KOTERET

DORON ZEILBERGER^{∗∗}

Section1

References

[1] Philip W. Anderson. What is wrong with QMC?, talk delivered at the 103rd Statistical Mechanics Conference,

    May 10, 2010, Rutgers University, http://www.math.rutgers.edu/events/smm/smm103-invitedtalks.html.

Doron Zeilberger, Mathematics Department, Rutgers University (New Brunswick), Piscataway, NJ, USA.

E-mail address: zeilberg@math.rutgers.edu

Date: May 26, 2010.

∗∗ Supported in part by the United States of America National Science Foundation.